# Autonomous Driving in Unstructured Environments: How Far Have We Come?

### v3

@@ III-B2 Vision-based Methods
TABLE VI: Vision-based traversability estimation methods for unstructured environments.
@@
Gao et al. [18] AlexNet Contrastive Learning Private data
Gao et al. [151] AlexNet Active and Contrastive Learning Private data, DeepScene
Sun et al. [152] PSPNet Supervised Learning ORFD
Seo et al. [153] PSPNet Self-Supervised Learning Private data
Jung et al. [19] ViT-H SAM Self-Supervised Learning RELLIS-3D
Chung et al. [154] EfficientNet-B0 Supervised Learning Private data
@@
Although LiDAR can accurately capture spatial structural information, its high cost has led to the proposal of many vision-based traversability estimation algorithms [155, 156]. Table VI lists the specific details of these methods.
@@
Sun et al. [152] enhance passable area extraction using a modified PSPNet with MobileNetv2 and dilation convolution modules, improving feature extraction and fusion. The method shows significant improvements in detecting passable areas in off-road environments. Chung et al. [154] propose a learning-based method for real-time long-range terrain elevation map prediction from onboard egocentric images. It includes a transformer-based encoder, orientation-aware positional encoding, and a history-augmented map embedding to enhance elevation map consistency and performance.
@@
Some algorithms employ contrastive learning by comparing traversable and non-traversable regions. Gao et al. [18] introduce a contrastive learning method for terrain analysis using human-annotated anchor patches to distinguish traversability levels, and develops fine-grained semantic segmentation and mapping techniques. Experimental validation shows effectiveness across varied off-road scenarios. Gao et al. [151] propose an active contrastive learning approach using patch-based weak annotations instead of pixelwise labels. It employs contrastive learning for feature representation and adaptive clustering for category discovery, with a risk evaluation technique for prioritizing high-risk frames for labeling.
@@
Some algorithms also adopt a self-supervised approach to reduce the reliance on data labeling. Seo et al. [153] present a self-supervised approach for learning traversability from images, deriving self-supervised labels from historical driving data and training a neural network with a one-class classification algorithm, integrating advanced self-supervised techniques. Jung et al. [19] introduce a self-supervised learning method for image-based traversability prediction using contrastive representation learning and instance-based segmentation masks, outperforming recent methods and showing compatibility with a model-predictive controller.
III-B3 Fusion-based Methods
@@
Some methods combine LiDAR’s geometric structure with camera texture information, using multi-sensor fusion for traversability estimation [157, 158]. Table VII presents the experimental results of fusion-based traversability estimation on the ORFD [52] dataset.
@@
TABLE VII: Quantitative results of fusion-based traversability estimation methods on the testing set of ORFD [52].
@@
FuseNet [159] $\checkmark$ $\checkmark$ 87.4% 74.5% 85.2% 79.5% 66.0% 50.0M 49.0 Hz
SNE-RoadSeg [160] $\checkmark$ $\checkmark$ 93.8% 86.7% 92.7% 89.6% 81.2% 201.3M 12.5 Hz
OFF-Net [52] $\checkmark$ $\checkmark$ 94.5% 86.6% 94.3% 90.3% 82.3% 25.2M 33.9 Hz
FSN-Swin [20] $\checkmark$ $\checkmark$ 96.4% 91.2% 95.8% 93.4% 87.7% - -
M2F2-Net [21] $\checkmark$ $\checkmark$ 98.1% 97.3% 95.5% 96.4% 93.1% 26.8M 34.0 Hz
OFF-CSUNet [161] $\checkmark$ $\checkmark$ 98.6% 97.7% 95.6% 96.9% 94.1% 31.1M 26.0Hz
NAIFNet [162] $\checkmark$ $\checkmark$ 98.4% 97.5% 96.4% 97.0% 94.1% 26.8M 11.5 Hz
Guan et al. [131] introduce the terrain traversability mapping and navigation system for autonomous excavators in unstructured environments, integrating RGB images and 3D point clouds into a global map for real-time navigation. VINet [163] is a visual and inertial-based network for terrain classification in robotic navigation, using a novel labeling scheme for generalization on unfamiliar surfaces and an adaptive control framework for improved navigation performance. Bae et al. [164] present a deep metric learning approach that integrates unlabeled data with few prototypes for binary segmentation and traversability regression, introducing a new evaluation metric for comprehensive performance assessment. Feng et al. [165] address untrustworthy features in depth images with the adaptive-mask fusion network, which uses adaptive-weight masks to merge RGB and depth images, improving segmentation accuracy.
@@
VPrs-Net [166] is a multi-task network for vanishing point detection and road segmentation, enhancing accuracy in off-road environments with VP-guided models and a multi-attention architecture. Evaluation on the ORFD dataset demonstrates its effectiveness. M2F2-Net [21] and OFF-CSUNet [161] are multi-modal networks for free-space detection, using multi-modal cross fusion and a multi-modal segmentation decoder to integrate RGB and surface normal maps, including an edge segmentation decoder module for improved road edge extraction. Duraisamy and Natarajan [167] introduce a framework for classifying and clustering regions, incorporating LiDAR-based ground segmentation with adaptive likelihood estimation and late fusion of deep learning and LiDAR results for improved drivable region classification. Xu et al. [168] combine terrain features from 3D point clouds and RGB images to create a global terrain traversal cost map, using Bayesian generalized kernel inference and Kalman filtering for real-time local elevation maps and probabilistic semantic mapping.
@@
III-B4 Semantic Segmentation
TABLE VIII: Learning-based semantic segmentation methods for unstructured environments.
Reference Architecture Sensor Dataset
Viswanath et al. [171] BiSeNet v2, HRNet v2+OCR RGB RELLIS-3D [25], RUGD [57]
Singh et al. [24] ViT, ResNet50, DINO RGB RELLIS-3D [25], RUGD [57]
Liu et al. [22] Euclidean Clustering, RANSAC LiDAR Gazebo [172]
Zhou et al. [23] PointTensor, Cylinder LiDAR RELLIS-3D [25], CARLA [173]
Viswanath et al. [174] FCN LiDAR RELLIS-3D [25]
Jiang and Saripalli [175] DeepLab v3 RADAR RELLIS-3D [25]
Lian et al. [176] Patchwork++, LIDROR, YOLO v5 RGB, LiDAR KITTI [177], Off-road
Feng et al. [26] RangeNet++, ResNet RGB, LiDAR RELLIS-3D [177], CARLA [173]
Kim and Seo [27] BKI, EDL RGB, LiDAR RELLIS-3D [25], Off-road
Li et al. [178] GPT-3.5, CLIP RGB IDD [179], ORFD [52], AutoMine [180]
Compared to traversability estimation, semantic segmentation offers a more comprehensive understanding of a vehicle’s surroundings [181], as shown in Fig. 7. With the advancement of deep learning, several semantic segmentation algorithms tailored for unstructured scenarios in autonomous driving have been proposed, drawing inspiration from mature semantic segmentation methods [182, 183, 184, 185]. Detailed information on these methods is provided in Table VIII.
LiDAR-based Methods
Vision-based Methods
Fusion-based Methods
@@ IV Path Planning
Path planning [187] and trajectory planning [188] algorithms are processes that create the optimal route to successfully reach the destination through navigable areas [189]. This optimality can be reflected in path length, energy consumption, or any other metrics deemed important by the operator [190]. Path planning methods are divided into global path planning and local path planning (Fig. 8). Because of the diversity and complexity of actual application scenarios, the requirements for path planning vary in different scenarios, so there are many types of path planning methods, as Table IX shows.
TABLE IX: Path Planning Methods.
Path Planning Type Advantage Disadvantage
Global Path Planning Search-based Excellent performance in handling static environments, small-scale maps, and problems with high requirements for optimality Computational efficiency is limited when dealing with high-dimensional spaces and large-scale environments
@@
Local Path Planning Optimization-based Can precisely model for specific optimization goals and constraints, and provide accurate solutions when considering specific goals May have higher computational complexity
@@
Compared to structured environments, path planning in unstructured environments faces several major challenges. These challenges include high uncertainty and dynamic changes, complex state spaces, limited prior knowledge, real-time operation requirements, and the complexity of long-term planning [191, 192]. This section will discuss these challenges in detail and explore path planning algorithms for unstructured environments.
IV-A Challenges of Path Planning in Unstructured Environment
IV-A1 Challenges of Global Path Planning
@@
IV-A2 Challenges of Local Path Planning
@@
IV-B Global Path Planning
IV-B1 Search-based Methods
The main idea behind search-based path planning algorithms [198] is to preprocess the planning space into a grid map containing obstacle information. In the grid map, nodes represent discrete states in the state space, and state transitions are achieved through connections between nodes. Compared to other path planning algorithms, search-based algorithms have higher solution stability, consistently producing the same planning results in identical scenarios. However, due to the discretization of the continuous planning space, the generated paths may not be smooth enough, introducing quantization errors. Table X presents the search-based path planning algorithms for unstructured environments.
TABLE X: Search-based path planning methods for unstructured environments.
@@ IV-B1 Search-based Methods
Stentz [28] Dynamic A*
Likhachev et al. [199] Anytime Dynamic A*
Montemerlo et al. [200] AD*
Ziegler et al. [201] A*, Voronoi Cost Function
Li et al. [202] Anytime Repairing A*
Roy et al. [203] Dynamic Programming
Sedighi et al. [204] Hybrid A*
Sakayori and Ishigami [205] Dijkstra’s Algorithm
Hong et al. [206] Improved A*
Hua et al. [207] PRM, Improved A*
Xu et al. [29] Velocity Hybrid A*
Dijkstra’s algorithm, an early search algorithm, is mainly used to find the shortest path between any two nodes in a connected graph. It adopts a breadth-first search approach, starting from the initial node and progressively evaluating nodes adjacent to the current node. Its search strategy is greedy, defining a cost function (usually the distance to the starting point) and selecting the node with the lowest cost each time to include in the path. If a feasible path exists between the start and end points, this algorithm guarantees finding the shortest feasible path. Sakayori and Ishigami [205] considere the power consumption model of planetary rovers on rugged terrain, designed a cost function accounting for both generation and consumption, and used Dijkstra’s algorithm to obtain path planning results. However, as a breadth-first search method, when the search range needs to be expanded or finer gridding is required, the number of nodes in the map will increase significantly, leading to a substantial increase in the total search volume of the algorithm and a rapid growth in the time to solve the optimal path. Therefore, Dijkstra’s algorithm is rarely directly used for large-scale or high-precision path planning problems.
Dynamic Programming (DP) solves optimal solutions by breaking problems into smaller subproblems. Its main idea is to utilize the overlapping nature of subproblems, reducing computational redundancy through memorization or recursion, significantly improving computational efficiency. Roy et al. [203] propose an efficient hierarchical method for vehicle path planning in vast off-road terrains. Global path planning uses the dynamic programming algorithm based on available low-resolution terrain information before the mission to find the optimal path outline. Subsequently, the local optimization program fine-tunes the optimal predefined route based on the results of the global program and integrates high-resolution terrain maps and obstacle detection algorithms.
@@
Improved versions of the A* algorithm include Dynamic A* [28], Anytime Repairing A*[202], Anytime Dynamic A* [199], and Hybrid A* [204]. In the 2007 DARPA Urban Challenge, the vehicle Boss, using the AD* algorithm, achieved great success, demonstrating its practical application [200]. The A* algorithm improves search efficiency through heuristic functions, performing exceptionally well in large-scale graph searches. Hong et al. [206] propose an improved A* algorithm based on terrain data for long-distance path planning tasks. This method uses terrain data maps generated from digital elevation models and optimizes the algorithm in data structure and retrieval strategy. As long as the heuristic function is reasonable (i.e., never overestimates the actual cost), the A* algorithm guarantees finding the optimal path. The heuristic function can be adjusted according to specific applications to meet different path planning needs.
Ziegler et al. [201] implement the A* algorithm combitned with a Voronoi cost function for planning in unstructured spaces and parking lots. However, in complex or large-scale search spaces, the A* algorithm may face high time and space complexity, consuming significant computational resources. Additionally, selecting an appropriate heuristic function is crucial for the algorithm’s performance, as an unsuitable choice may lead to inefficient searches. Stentz [28] proposes a new algorithm, D*, capable of planning paths efficiently, optimally, and completely in unknown, partially known, and changing environments.
Hua et al. [207] propose a path planning method considering terrain factors and soil mechanics’ impact on UGV maneuverability. They used the probabilistic roadmap method to establish a connection matrix and a multidimensional traffic cost evaluation matrix between sampling nodes and proposed an improved A* algorithm based on vehicle mobility costs to generate global paths. Xu et al. [29] propose a centralized decision-making distributed planning framework to solve the Multi-Vehicle Trajectory Planning (MVTP) problem in unstructured conflict areas, using Velocity Hybrid A* (V-Hybrid A*) for trajectory search.
IV-B2 Sample-based Methods
Sampling-based path planning algorithms typically have probabilistic completeness, meaning they can find the optimal path that exists on the planning map given a sufficient number of samples. These methods introduce randomness into the path planning process, with the most common including the Probabilistic Roadmap (PRM) [208] and the RRT methods. Table XI lists the sample-based path planning algorithms for unstructured environments.
TABLE XI: Sample-based path planning methods for unstructured environments.

### v4

@@ 3.2.2 Semantic Segmentation
Table 6: Learning-based semantic segmentation methods for unstructured environments.
Reference Architecture Sensor Dataset
[196] BiSeNet v2, HRNet v2+OCR RGB RELLIS-3D  RUGD
[177] ViT, ResNet50, DINO RGB RELLIS-3D, RUGD
[114] Euclidean Clustering, RANSAC LiDAR Gazebo
[253] PointTensor, Cylinder LiDAR RELLIS-3D, CARLA
[195] FCN LiDAR RELLIS-3D
[79] DeepLab v3 RADAR RELLIS-3D
[107] Patchwork++, LIDROR, YOLO v5 RGB, LiDAR KITTI, Off-road
[38] RangeNet++, ResNet RGB, LiDAR RELLIS-3D, CARLA
[87] BKI, EDL RGB, LiDAR RELLIS-3D, Off-road
[106] GPT-3.5, CLIP RGB IDD, ORFD, AutoMine
Compared to traversability estimation, semantic segmentation offers a more comprehensive understanding of a vehicle’s surroundings [239], as shown in Fig. 6. With the advancement of deep learning, several semantic segmentation algorithms tailored for unstructured scenarios in autonomous driving have been proposed, drawing inspiration from mature semantic segmentation methods [159, 20, 250, 252]. Detailed information on these methods is provided in Table 6.
3.3 RGB-based Methods
3.3.1 Traversability Estimation
Table 7: Vision-based traversability estimation methods for unstructured environments.
@@ 3.3.1 Traversability Estimation
[46] AlexNet Contrastive Learning Private data
[47] AlexNet Active and Contrastive Learning Private data, DeepScene
[181] PSPNet Supervised Learning ORFD
[167] PSPNet Self-Supervised Learning Private data
[80] ViT-H SAM Self-Supervised Learning RELLIS-3D
[23] EfficientNet-B0 Supervised Learning Private data
@@
Although LiDAR can accurately capture spatial structural information, its high cost has led to the proposal of many vision-based traversability estimation algorithms [221, 202]. Table 7 lists the specific details of these methods.
@@
[181] enhance passable area extraction using a modified PSPNet with MobileNetv2 and dilation convolution modules, improving feature extraction and fusion. The method shows significant improvements in detecting passable areas in off-road environments. [23] propose a learning-based method for real-time long-range terrain elevation map prediction from onboard egocentric images. It includes a transformer-based encoder, orientation-aware positional encoding, and a history-augmented map embedding to enhance elevation map consistency and performance.
@@
Some algorithms employ contrastive learning by comparing traversable and non-traversable regions. [46] introduce a contrastive learning method for terrain analysis using human-annotated anchor patches to distinguish traversability levels, and develops fine-grained semantic segmentation and mapping techniques. Experimental validation shows effectiveness across varied off-road scenarios. [47] propose an active contrastive learning approach using patch-based weak annotations instead of pixelwise labels. It employs contrastive learning for feature representation and adaptive clustering for category discovery, with a risk evaluation technique for prioritizing high-risk frames for labeling.
@@
Some algorithms also adopt a self-supervised approach to reduce the reliance on data labeling. [167] present a self-supervised approach for learning traversability from images, deriving self-supervised labels from historical driving data and training a neural network with a one-class classification algorithm, integrating advanced self-supervised techniques. [80] introduce a self-supervised learning method for image-based traversability prediction using contrastive representation learning and instance-based segmentation masks, outperforming recent methods and showing compatibility with a model-predictive controller.
3.3.2 Semantic Segmentation
3.4 RADAR-based Methods
In unstructured off-road environments, perception systems often need to cope with harsh conditions such as dust, fog, rain, snow, and vegetation occlusion, which can severely degrade the performance of conventional vision and LiDAR sensors. RADAR, due to its strong penetration capability and environmental robustness, has gradually become an important perception modality for off-road autonomous driving. Unlike sensors that rely on visible light or laser reflection, RADAR can detect obstacles and estimate relative velocity in low-visibility or occluded scenarios, making it particularly suitable for deserts, farmland, or post-disaster environments. In recent years, millimeter-wave and frequency-modulated continuous-wave (FMCW) RADAR have attracted considerable attention for 3D environment perception and terrain reconstruction. For example, This paper demonstrates that RADAR, adapted from existing LiDAR-based methods, provides longer-range perception, robust ground plane estimation and obstacle detection, and enables successful autonomous navigation in challenging off-road environments.
3.5 Thermal-based Methods
Thermal imaging plays an important role in off-road autonomous driving. Compared with traditional RGB vision, thermal sensors can perceive the environment reliably under challenging conditions such as nighttime, glare, dust, fog, or vegetation occlusion. They are independent of visible light, highly resistant to interference, and capable of detecting living beings, obstacles, and terrain boundaries, thereby enhancing safety and robustness in extreme or low-visibility environments. However, thermal images contain limited texture and lower resolution, making them insufficient for fine-grained perception alone. Therefore, they are often fused with RGB or LiDAR data to leverage complementary information across modalities, significantly improving environmental understanding and navigation reliability in off-road scenarios.
3.6 Fusion-based Methods
3.6.1 Traversability Estimation
@@
Some methods combine LiDAR’s geometric structure with camera texture information, using multi-sensor fusion for traversability estimation [222, 220]. Table 8 presents the experimental results of fusion-based traversability estimation on the ORFD [125] dataset.
@@
Table 8: Quantitative results of fusion-based traversability estimation methods on the testing set of ORFD [125].
@@
FuseNet $\checkmark$ $\checkmark$ 87.4% 74.5% 85.2% 79.5% 66.0% 50.0M 49.0 Hz
SNE-RoadSeg $\checkmark$ $\checkmark$ 93.8% 86.7% 92.7% 89.6% 81.2% 201.3M 12.5 Hz
OFF-Net $\checkmark$ $\checkmark$ 94.5% 86.6% 94.3% 90.3% 82.3% 25.2M 33.9 Hz
FSN-Swin $\checkmark$ $\checkmark$ 96.4% 91.2% 95.8% 93.4% 87.7% - -
M2F2-Net $\checkmark$ $\checkmark$ 98.1% 97.3% 95.5% 96.4% 93.1% 26.8M 34.0 Hz
OFF-CSUNet $\checkmark$ $\checkmark$ 98.6% 97.7% 95.6% 96.9% 94.1% 31.1M 26.0Hz
NAIFNet $\checkmark$ $\checkmark$ 98.4% 97.5% 96.4% 97.0% 94.1% 26.8M 11.5 Hz
[55] introduce the terrain traversability mapping and navigation system for autonomous excavators in unstructured environments, integrating RGB images and 3D point clouds into a global map for real-time navigation. VINet [56] is a visual and inertial-based network for terrain classification in robotic navigation, using a novel labeling scheme for generalization on unfamiliar surfaces and an adaptive control framework for improved navigation performance. [5] present a deep metric learning approach that integrates unlabeled data with few prototypes for binary segmentation and traversability regression, introducing a new evaluation metric for comprehensive performance assessment. [39] address untrustworthy features in depth images with the adaptive-mask fusion network, which uses adaptive-weight masks to merge RGB and depth images, improving segmentation accuracy.
@@
VPrs-Net [116] is a multi-task network for vanishing point detection and road segmentation, enhancing accuracy in off-road environments with VP-guided models and a multi-attention architecture. Evaluation on the ORFD dataset demonstrates its effectiveness. M2F2-Net [237] and OFF-CSUNet [154] are multi-modal networks for free-space detection, using multi-modal cross fusion and a multi-modal segmentation decoder to integrate RGB and surface normal maps, including an edge segmentation decoder module for improved road edge extraction. [32] introduce a framework for classifying and clustering regions, incorporating LiDAR-based ground segmentation with adaptive likelihood estimation and late fusion of deep learning and LiDAR results for improved drivable region classification. [225] combine terrain features from 3D point clouds and RGB images to create a global terrain traversal cost map, using Bayesian generalized kernel inference and Kalman filtering for real-time local elevation maps and probabilistic semantic mapping.
@@
3.6.2 Semantic Segmentation
@@ 4 Path Planning
Path planning [49] and trajectory planning [173] algorithms are processes that create the optimal route to successfully reach the destination through navigable areas [201]. This optimality can be reflected in path length, energy consumption, or any other metrics deemed important by the operator [7]. Path planning methods are divided into global path planning and local path planning (Fig. 7). Because of the diversity and complexity of actual application scenarios, the requirements for path planning vary in different scenarios, so there are many types of path planning methods, as Table 9 shows.
Table 9: The advantage and disadvantage of different path planning methods.
Type Advantage Disadvantage
Global Search-based Excellent performance in handling static environments, small-scale maps, and problems with high requirements for optimality Computational efficiency is limited when dealing with high-dimensional spaces and large-scale environments
@@
Local Optimization-based Can precisely model for specific optimization goals and constraints, and provide accurate solutions when considering specific goals May have higher computational complexity
@@
Compared to structured environments, path planning in unstructured environments faces several major challenges. These challenges include high uncertainty and dynamic changes, complex state spaces, limited prior knowledge, real-time operation requirements, and the complexity of long-term planning [22, 36, 59]. This section will discuss these challenges in detail and explore path planning algorithms for unstructured environments.
4.1 Challenges of Path Planning in Unstructured Environment
4.1.1 Challenges of Global Path Planning
@@
4.1.2 Challenges of Local Path Planning
@@
4.2 Global Path Planning
4.2.1 Search-based Methods
The main idea behind search-based path planning algorithms [163] is to preprocess the planning space into a grid map containing obstacle information. In the grid map, nodes represent discrete states in the state space, and state transitions are achieved through connections between nodes. Compared to other path planning algorithms, search-based algorithms have higher solution stability, consistently producing the same planning results in identical scenarios. However, due to the discretization of the continuous planning space, the generated paths may not be smooth enough, introducing quantization errors. Table 10 presents the search-based path planning algorithms for unstructured environments.
Table 10: Search-based path planning methods for unstructured environments.
@@ 4.2.1 Search-based Methods
[179] Dynamic A*
[109] Anytime Dynamic A*
[132] AD*
[258] A*, Voronoi Cost Function
[101] Anytime Repairing A*
[160] Dynamic Programming
[164] Hybrid A*
[162] Dijkstra’s Algorithm
[66] Improved A*
[72] PRM, Improved A*
[224] Velocity Hybrid A*
Dijkstra’s algorithm, an early search algorithm, is mainly used to find the shortest path between any two nodes in a connected graph. It adopts a breadth-first search approach, starting from the initial node and progressively evaluating nodes adjacent to the current node. Its search strategy is greedy, defining a cost function (usually the distance to the starting point) and selecting the node with the lowest cost each time to include in the path. If a feasible path exists between the start and end points, this algorithm guarantees finding the shortest feasible path. [162] considere the power consumption model of planetary rovers on rugged terrain, designed a cost function accounting for both generation and consumption, and used Dijkstra’s algorithm to obtain path planning results. However, as a breadth-first search method, when the search range needs to be expanded or finer gridding is required, the number of nodes in the map will increase significantly, leading to a substantial increase in the total search volume of the algorithm and a rapid growth in the time to solve the optimal path. Therefore, Dijkstra’s algorithm is rarely directly used for large-scale or high-precision path planning problems.
Dynamic Programming (DP) solves optimal solutions by breaking problems into smaller subproblems. Its main idea is to utilize the overlapping nature of subproblems, reducing computational redundancy through memorization or recursion, significantly improving computational efficiency. [160] propose an efficient hierarchical method for vehicle path planning in vast off-road terrains. Global path planning uses the dynamic programming algorithm based on available low-resolution terrain information before the mission to find the optimal path outline. Subsequently, the local optimization program fine-tunes the optimal predefined route based on the results of the global program and integrates high-resolution terrain maps and obstacle detection algorithms.
@@
Improved versions of the A* algorithm include Dynamic A* [179], Anytime Repairing A*[101], Anytime Dynamic A* [109], and Hybrid A* [164]. In the 2007 DARPA Urban Challenge, the vehicle Boss, using the AD* algorithm, achieved great success, demonstrating its practical application [132]. The A* algorithm improves search efficiency through heuristic functions, performing exceptionally well in large-scale graph searches. [66] propose an improved A* algorithm based on terrain data for long-distance path planning tasks. This method uses terrain data maps generated from digital elevation models and optimizes the algorithm in data structure and retrieval strategy. As long as the heuristic function is reasonable (i.e., never overestimates the actual cost), the A* algorithm guarantees finding the optimal path. The heuristic function can be adjusted according to specific applications to meet different path planning needs.
[258] implement the A* algorithm combitned with a Voronoi cost function for planning in unstructured spaces and parking lots. However, in complex or large-scale search spaces, the A* algorithm may face high time and space complexity, consuming significant computational resources. Additionally, selecting an appropriate heuristic function is crucial for the algorithm’s performance, as an unsuitable choice may lead to inefficient searches. [179] proposes a new algorithm, D*, capable of planning paths efficiently, optimally, and completely in unknown, partially known, and changing environments.
[72] propose a path planning method considering terrain factors and soil mechanics’ impact on UGV maneuverability. They used the probabilistic roadmap method to establish a connection matrix and a multidimensional traffic cost evaluation matrix between sampling nodes and proposed an improved A* algorithm based on vehicle mobility costs to generate global paths. [224] propose a centralized decision-making distributed planning framework to solve the Multi-Vehicle Trajectory Planning (MVTP) problem in unstructured conflict areas, using Velocity Hybrid A* (V-Hybrid A*) for trajectory search.
4.2.2 Sample-based Methods
Sampling-based path planning algorithms typically have probabilistic completeness, meaning they can find the optimal path that exists on the planning map given a sufficient number of samples. These methods introduce randomness into the path planning process, with the most common including the Probabilistic Roadmap (PRM) [83] and the RRT methods. Table 11 lists the sample-based path planning algorithms for unstructured environments.
Table 11: Sample-based path planning methods for unstructured environments.